emotional quotient (EQ) score on the least version of Sirug EQ Test are transformed so as follow a normal distribution with a mean of 110 and standard deviation of 15. find the 96th percentile of the distribution of Emotional Quotient

1
Expert's answer
2021-11-18T06:49:41-0500

"x=\mu+Z\sigma"

Firstly needs to determine Z value using normal distribution table and 0.96;

"x=110+1.75\times15"

"x=110+26.25=136.25"

The means 96th percentile of the distribution of Emotional Quotient is 136.25.